Equipment acquired on January 8 at a cost of $168,000 has an estimated useful life of 18 years, has an estimated residual value of $15,000, and is depreciated by the straight-line method.

A. What was the book value of the equipment at December 31 the end of the fourth year?
B. Assuming that the equipment was sold on April 1 of the fifth year for $125,000, journalize the entries to record (1) depreciation for the three months until the sale date and (2) the sale of the equipment.

X

First Question

A. What was the book value of the equipment at December 31 the end of the fourth year?

. Assuming that the equipment was sold on April 1 of the fifth year for $125,000, journalize the entries to record the following (refer to the Chart of Accounts for exact wording of account titles):

1. Depreciation for the three months until the sale date

2. The sale of the equipment
Answer #1

Cost of machinery = $168,000

Residual value = $15,000

Useful life = 18 years

Annual depreciation = (Cost of machinery-Residual value)/ useful life

= (168,000-15,000)/18

= $8,500

Accumulated depreciation for 4 years = Annual depreciation x 4

= 8,500 x 4

= $34,000

A.

Book value of equipment at December 31 at the end of year 4 = Cost of equipment - Accumulated depreciation for 4 years

= 168,000-34,000

= $134,000

Depreciation for 3 months of year 5 = Annual depreciation x 3/12

= 8,500 x 3/12

= $2,125

Accumulated depreciation at April 1, year 5 = Accumulated depreciation for 4 years + Depreciation for 3 months of year 5

= 34,000+2,125

= $36,125

Sale of equipment = $125,000

Book value of equipment at April 1, year 5 = Cost of equipment - Accumulated depreciation at April 1, year 5

= 168,000-36,125

= $131,875

Loss on sale of equipment = Book value - Sale of equipment

= 131,875-125,000

= $6,875

Date General Journal Debit Credit
Year 5, April 1 Depreciation expense $2,125
Accumulated depreciation- equipment $2,125
( To record depreciation expense)
Year 5, April 1 Cash $125,000
Loss on sale of equipment $6,875
Accumulated depreciation- equipment $36,125
Equipment $168,000
( To record sale of the equipment)

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks